# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v6

@@ Abstract
Index Terms:
text-to-SQL, large language models, database, natural language understanding
@@ I Introduction
Figure 1: An example of LLM-based text-to-SQL is selected from the BIRD [1] dataset. A user asks a question about football leagues. The LLM takes this question along with the schema of the corresponding database as input and generates an SQL query as output. The generated SQL query can be executed in the database, retrieves the content “The 5 leagues with the highest matches”, providing the answer to the user’s question.
@@
Text-to-SQL is a long-standing task in natural language processing, aiming to convert natural language (NL) questions into database-executable SQL queries. Fig. 1 illustrates an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input to generate an SQL query. Executing this SQL query in the database retrieves the relevant content needed to answer the user’s question. This system serves as a natural language interface to the database (NLIDB) using LLMs. According to the 2023 Stack Overflow survey, SQL is one of the most widely used programming languages, with over half of professional developers (51.52%) employing it, while only about a third (35.29%) having received systematic training. The NLIDB therefore allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5]. The significant value of text-to-SQL has triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
@@
Previous traditional methods have made notable progress in implementing text-to-SQL. As shown in Fig. 2, these implementations have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [9], specifically suitable for simple database scenarios. In recent years, due to the heavy labor costs [10] associated with rule-based methods and the growing complexity of database environments [11, 12, 13], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [14, 15], which can automatically learn a mapping from the user question to its corresponding SQL [16, 17]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capabilities have become the new paradigm for text-to-SQL systems [18], boosting their performance to a new level [19, 20, 21]. Incremental research on PLM-based optimization, such as table content encoding [18, 17, 22] and pre-training [18, 23], has further advanced this field. Most recently, LLM-based approaches have implemented text-to-SQL through the in-context learning [7] and fine-tuning [24] paradigms, reaching state-of-the-art accuracy with well-designed frameworks and stronger comprehension capabilities compared to PLM-based methods.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by a rich training corpus [26, 27]. Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ LLM-based Text-to-SQL, leaf-head, ver [ Datasets
III-A), leaf-datasets, text width=5em [ Original
Datasets
Post-annotated
Datasets, leaf-datasets, text width=6em [ Cross-
domain, leaf-datasets, text width=5.5em [BIRD [1], DuSQL [31], CoSQL [32], Spider [11], WikiSQL [12], KaggleDBQA [33], ADVETA [34],
Spider-SS [35], Spider-CG [35], Spider-DK [36], Spider-SYN [37], Spider-Realistic [38], CSpider [39],
SParC [40], modelnode-datasets, text width=43.5em] ] [ Knowledge-
augmented, leaf-datasets, text width=5.5em [BIRD [1], SQUALL [41], Spider-DK [36], BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Cross-
lingual, leaf-datasets, text width=5.5em [DuSQL [31], CSpider [39], Spider-Vietnamese [43], modelnode-datasets, text width=43.5em] ] [ Context-
dependent, leaf-datasets, text width=5.5em [CoSQL [32], Spider-SS [35], Spider-CG [35], SparC [40], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [34], Spider-SYN [37], Spider-Realistic [38], Dr. Spider [44], modelnode-datasets, text width=43.5em] ] [ Long-context, leaf-datasets, text width=5.5em [BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Specialized-
domain, leaf-datasets, text width=5.5em [BULL [45], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
Metrics
III-B), leaf-metrics,text width=5em [ Content-
Matching
based, leaf-metrics, text width=6em [Component Matching (CM) [11], Exact Matching (EM) [11], modelnode-metrics, text width=50.7em] ] [ Execution
based, leaf-metrics, text width=6em [Execution Accuracy (EX) [11], Valid Efficiency Score (VES) [1], modelnode-metrics, text width=50.7em] ] ] [ Methods
IV), leaf-methods,text width=5em [ In-context
Learning
Paradigm
IV-A), leaf-methods, text width=6em [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [46], DIN-SQL [7], QDecomp [47], C3 [28], MAC-SQL [48], DEA-SQL [49],
SGU-SQL [30], MetaSQL [50], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], TA-SQL [54], MCS-SQL [55],
CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], Spider-Agent [42],
ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Prompt
Optimization, leaf-methods, text width=6.5em [DESEM+P [63], StructGPT [64], SD+SA+Voting [65], RAG+SP&DRC [66], C3 [28], DAIL-SQL [8],
ODIS [67], ACT-SQL [68], FUSED [69], DELLM [70], Dubo-SQL [71], TA-SQL [54],
MCS-SQL [55], SuperSQL [57], Distillery [58], RSL-SQL [59], ICRL [72], Gen-SQL [60],
SAFE-SQL [73], ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Reasoning
Enhancement, leaf-methods, text width=6.5em[CoT [47, 1, 8, 30], QDecomp [47], Least-to-Most [47], SQL-PaLM [74], ACT-SQL [68], POT [75],
SQL-CRAFT [75], FUXI [76], SuperSQL [57], CHASE-SQL [77], ReFoRCE [61], Spider-Agent [42], modelnode-methods,text width=42.5em] ] [ Execution
Refinement, leaf-methods, text width=6.5em[MBR-Exec [78], Coder-Reviewer [46], LEVER [79], SELF-DEBUGGING [80], DESEM+P [63],
DIN-SQL [7], SD+SA+Voting [65], SQL-PaLM [74], RAG+SP&DRC [66], C3 [28], MAC-SQL [48],
DELLM [70], SQL-CRAFT [75], FUXI [76], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], Dubo-SQL [71],
MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [81], CHASE-SQL [77], RSL-SQL [59],
Spider-Agent [42], ReFoRCE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
Paradigm
IV-B), leaf-methods,text width=6em [ Enhanced
Architecture, leaf-methods, text width=6.5em [CLLMs [82], modelnode-methods,text width=42.5em] ] [ Pre-training, leaf-methods, text width=6.5em [CodeS [24], modelnode-methods,text width=42.5em] ] [ Data
Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [8], Symbol-LLM [83], CodeS [24], StructLM [84], Dubo-SQL [71], Distillery [58],
XiYan-SQL [85], SHARE [86], modelnode-methods,text width=42.5em] ] [ Multi-task
Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [87], DTS-SQL [88], KaSLA [29], MSc-SQL [89], XiYan-SQL [85], ROUTE [90],
SHARE [86], modelnode-methods,text width=42.5em] ] ] ] ]
@@
Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [91]; nodes are ordered by release time.
@@
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [92]. Recent related works have surveyed text-to-SQL systems using deep learning approaches and provided insights into previous deep neural network and PLM-based research [93, 3, 27, 94]. In this survey, we aim to catch up with recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of topics related to LLM-based text-to-SQL, including:
@@ II Overview
Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database. It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [95]. This capability can greatly benefit domains such as business intelligence, customer support, and scientific research by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@ II-A1 Linguistic Complexity and Ambiguity
Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding parts of SQL queries [38]. Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [96, 97]. Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [1].
@@ II-A2 Schema Understanding and Representation
To generate accurate SQL queries, text-to-SQL systems must have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables. Representing and encoding schema information is challenging because the database schema is often complex and varies significantly across different domains [11].
@@ II-A4 Cross-Domain Generalization
Text-to-SQL systems often struggle to generalize across various database scenarios and domains. Models trained on a specific domain may not perform well on questions from other domains due to variations in vocabulary, database schema structure, and question patterns. Developing systems that can effectively generalize to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [98].
@@ II-B1 Rule-based Methods
Early text-to-SQL systems relied heavily on rule-based methods [9, 10, 23], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. Rule-based systems excel at generating syntactically correct SQL queries by relying on predefined templates and strict grammatical rules, ensuring adherence to SQL syntax and structure. Their deterministic nature guarantees consistent outputs for specific input patterns, minimizing variability in query generation. However, these methods struggle to interpret linguistically complex or ambiguous natural language questions, such as those involving nested clauses, coreferences, or ellipses, and often fail to map these elements to database schema components like tables or columns. Their reliance on manual rules limits generalization to unseen question patterns or schema structures, requiring frequent updates for new domains. Additionally, encoding complex schemas with multi-table relationships into usable rules is labor-intensive and prone to errors.

### v7

@@ I Introduction
Figure 1: An example of LLM-based text-to-SQL is selected from the BIRD [li2023BIRD] dataset. A user asks a question about football leagues. The LLM takes this question along with the schema of the corresponding database as input and generates an SQL query as output. The generated SQL query can be executed in the database, retrieves the content “The 5 leagues with the highest matches”, providing the answer to the user’s question.
@@
Text-to-SQL is a long-standing task in natural language processing, aiming to convert natural language (NL) questions into database-executable SQL queries. Fig. 1 illustrates an example of a large language model-based (LLM-based) text-to-SQL system. Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input to generate an SQL query. Executing this SQL query in the database retrieves the relevant content needed to answer the user’s question. This system serves as a natural language interface to the database (NLIDB) using LLMs. According to the 2023 Stack Overflow survey, SQL is one of the most widely used programming languages, with over half of professional developers (51.52%) employing it, while only about a third (35.29%) having received systematic training. The NLIDB therefore allows non-skilled users to access structured databases like professional database engineers [wang2022proton, qin2022survey] and enhances human-computer interaction [xu2020autoqa]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [lin2021truthfulqa]. The significant value of text-to-SQL has triggered a range of studies on its integration and optimization with LLMs [rajkumar2022evaluating, pourreza2023dinsql, gao2023dailsql]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
@@
Previous traditional methods have made notable progress in implementing text-to-SQL. As shown in Fig. 2, these implementations have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [li2014constructing], specifically suitable for simple database scenarios. In recent years, due to the heavy labor costs [mahmud2015rule] associated with rule-based methods and the growing complexity of database environments [yu2018Spider, zhong2017WikiSQL, pourreza2023evaluating], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [sutskever2014sequence, vaswani2017attention], which can automatically learn a mapping from the user question to its corresponding SQL [hui2021improving, choi2021ryansql]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capabilities have become the new paradigm for text-to-SQL systems [yin2020tabert], boosting their performance to a new level [li2023resdsql, li2023graphix, rai2023improving]. Incremental research on PLM-based optimization, such as table content encoding [yin2020tabert, choi2021ryansql, lyu2020hybrid] and pre-training [yin2020tabert, yu2021grappa], has further advanced this field. Most recently, LLM-based approaches have implemented text-to-SQL through the in-context learning [pourreza2023dinsql] and fine-tuning [li2024codes] paradigms, reaching state-of-the-art accuracy with well-designed frameworks and stronger comprehension capabilities compared to PLM-based methods.
@@
The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3. SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer. LLMs have proven to perform a successful vanilla implementation [rajkumar2022evaluating, liu2023comprehensive], benefiting from the powerful semantic parsing capacity enabled by a rich training corpus [yang2024harnessing, zhao2023survey]. Further studies on enhancing LLMs for question understanding [pourreza2023dinsql, gao2023dailsql], schema comprehension [dong2023c3, yuan2025knapsack], and SQL generation [zhang2024sgusql] are being increasingly released.
@@
[…]
LLM-based Text-to-SQL
Datasets III-A) Original Datasets & Post-annotated Datasets Cross- domain BIRD [li2023BIRD], DuSQL [wang2020DuSQL], CoSQL [yu2019CoSQL], Spider [yu2018Spider], WikiSQL [zhong2017WikiSQL], KaggleDBQA [lee2021KaggleDBQA], ADVETA [pi2022ADVETA], Spider-$S$ [gan2022Spider-CG], Spider-CG [gan2022Spider-CG], Spider-DK [gan2021Spider-DK], Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], CSpider [min2019CSpider], SParC [yu2019SParC] Knowledge- augmented BIRD [li2023BIRD], SQUALL [shi2020SQUALL], Spider-DK [gan2021Spider-DK], BIRD-CRITIC [li2023BIRD], Spider 2.0 [lei2025spider2] Cross- lingual DuSQL [wang2020DuSQL], CSpider [min2019CSpider], Spider-Vietnamese [tuan2020pilot] Context- dependent CoSQL [yu2019CoSQL], Spider-$S$ [gan2022Spider-CG], Spider-CG [gan2022Spider-CG], SparC [yu2019SParC] Robustness ADVETA [pi2022ADVETA], Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], Dr. Spider [chang2023dr] Long-context BIRD-CRITIC [li2023BIRD], Spider 2.0 [lei2025spider2] Specialized- domain BULL [zhang2024finsql] Evaluation Metrics III-B) Content- Matching based Component Matching (CM) [yu2018Spider], Exact Matching (EM) [yu2018Spider] Execution based Execution Accuracy (EX) [yu2018Spider], Valid Efficiency Score (VES) [li2023BIRD] Methods IV) In-context Learning Paradigm IV-A) Decomposition Coder-Reviewer [zhang2022coder], DIN-SQL [pourreza2023dinsql], QDecomp [tai2023exploring], C3 [dong2023c3], MAC-SQL [wang2024macsql], DEA-SQL [xie2024deasql], SGU-SQL [zhang2024sgusql], MetaSQL [fan2024metasql], $\textit{{R}}^{3}$ [xia2024r3], PET-SQL [li2024petsql], PURPLE [ren2024purple], TA-SQL [qu2024generation], MCS-SQL [lee2025mcs], CHESS [talaei2024chess], SuperSQL [li2024dawn], Distillery [maamari2024the], RSL-SQL [cao2024rsl], Gen-SQL [shi2025gen], Spider-Agent [lei2025spider2], ReFoRCE [deng2025reforce], LinkAlign [wang2025linkalign] Prompt Optimization DESEM+P [guo2023desem], StructGPT [jiang2023structgpt], SD+SA+Voting [nan2023enhancing], RAG+SP&DRC [guo2023ragsql], C3 [dong2023c3], DAIL-SQL [gao2023dailsql], ODIS [chang2023selective], ACT-SQL [zhang2023actsql], FUSED [wang2024fused], DELLM [hong2024knowledgetosql], Dubo-SQL [thorpe2024dubo], TA-SQL [qu2024generation], MCS-SQL [lee2025mcs], SuperSQL [li2024dawn], Distillery [maamari2024the], RSL-SQL [cao2024rsl], ICRL [toteja2025context], Gen-SQL [shi2025gen], SAFE-SQL [lee2025safe], ReFoRCE [deng2025reforce], LinkAlign [wang2025linkalign] Reasoning Enhancement CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql], QDecomp [tai2023exploring], Least-to-Most [tai2023exploring], SQL-PaLM [sun2023sqlpalm], ACT-SQL [zhang2023actsql], POT [xia2024sqlcraft], SQL-CRAFT [xia2024sqlcraft], FUXI [gu2024fuxi], SuperSQL [li2024dawn], CHASE-SQL [pourreza2025chasesql], ReFoRCE [deng2025reforce], Spider-Agent [lei2025spider2] Execution Refinement MBR-Exec [shi2022natural], Coder-Reviewer [zhang2022coder], LEVER [ni2023lever], SELF-DEBUGGING [chen2024selfdebugging], DESEM+P [guo2023desem], DIN-SQL [pourreza2023dinsql], SD+SA+Voting [nan2023enhancing], SQL-PaLM [sun2023sqlpalm], RAG+SP&DRC [guo2023ragsql], C3 [dong2023c3], MAC-SQL [wang2024macsql], DELLM [hong2024knowledgetosql], SQL-CRAFT [xia2024sqlcraft], FUXI [gu2024fuxi], $\textit{{R}}^{3}$ [xia2024r3], PET-SQL [li2024petsql], PURPLE [ren2024purple], Dubo-SQL [thorpe2024dubo], MCS-SQL [lee2025mcs], CHESS [talaei2024chess], Distillery [maamari2024the], E-SQL [caferouglu2024esql], CHASE-SQL [pourreza2025chasesql], RSL-SQL [cao2024rsl], Spider-Agent [lei2025spider2], ReFoRCE [deng2025reforce] Fine-tuning Paradigm IV-B) Enhanced Architecture CLLMs [kou2024cllms] Pre-training CodeS [li2024codes] Data Augmentation DAIL-SQL [gao2023dailsql], Symbol-LLM [xu2024symbolllm], CodeS [li2024codes], StructLM [zhuang2024structlm], Dubo-SQL [thorpe2024dubo], Distillery [maamari2024the], XiYan-SQL [gao2024xiyan], SHARE [qu2025share] Multi-task Tuning SQL-LLaMA [roziere2023codellama], DTS-SQL [pourreza2024dtssql], KaSLA [yuan2025knapsack], MSc-SQL [gorti2024mscsql], XiYan-SQL [gao2024xiyan], ROUTE [qin2025route], SHARE [qu2025share]
@@
Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [xu2023large]; nodes are ordered by release time.
@@
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [katsogiannis2023survey]. Recent related works have surveyed text-to-SQL systems using deep learning approaches and provided insights into previous deep neural network and PLM-based research [deng2022recent, qin2022survey, zhao2023survey, zhang2024tabluarsurvey]. In this survey, we aim to catch up with recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs. Specifically, the body of our survey covers a range of topics related to LLM-based text-to-SQL, including:
@@ II Overview
Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database. It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [ma2021mt]. This capability can greatly benefit domains such as business intelligence, customer support, and scientific research by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@ II-A1 Linguistic Complexity and Ambiguity
Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding parts of SQL queries [deng2021Spider-Realistic]. Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [rajpurkar2016squad, rajpurkar2018know]. Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [li2023BIRD].
@@ II-A2 Schema Understanding and Representation
To generate accurate SQL queries, text-to-SQL systems must have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables. Representing and encoding schema information is challenging because the database schema is often complex and varies significantly across different domains [yu2018Spider].
@@ II-A4 Cross-Domain Generalization
Text-to-SQL systems often struggle to generalize across various database scenarios and domains. Models trained on a specific domain may not perform well on questions from other domains due to variations in vocabulary, database schema structure, and question patterns. Developing systems that can effectively generalize to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [yang2024unveiling].
@@ II-B1 Rule-based Methods
Early text-to-SQL systems relied heavily on rule-based methods [li2014constructing, mahmud2015rule, yu2021grappa], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. Rule-based systems excel at generating syntactically correct SQL queries by relying on predefined templates and strict grammatical rules, ensuring adherence to SQL syntax and structure. Their deterministic nature guarantees consistent outputs for specific input patterns, minimizing variability in query generation. However, these methods struggle to interpret linguistically complex or ambiguous natural language questions, such as those involving nested clauses, coreferences, or ellipses, and often fail to map these elements to database schema components like tables or columns. Their reliance on manual rules limits generalization to unseen question patterns or schema structures, requiring frequent updates for new domains. Additionally, encoding complex schemas with multi-table relationships into usable rules is labor-intensive and prone to errors.